1. Suppose you have two random variables, X and Y with joint distribution given by the following tables So, for example, the probability that Y o,x - 0 is 4, and the probability that Y (a) Find the marginal distributions (pmfs) of X and Y, denoted f(x),J(Y). (b) Find the conditional distribution (pmf) of Y give X, denoted f(YX). (c) Find the expected values of X and Y, EX), E(Y). (d) Find the variances of X and Y, Var(X),Var(Y). (e) Find the covariance of X and Y, Cov(X, Y). (f) Find the correlation of X and Y, p(X, Y). (8) Are X and Y independent? Why or why not? 0+2-6
Answer #1

The joint distribution function of two random variables X and Y is given as follows: Y = 0 X=0X=1 0.4 0.2 0 .1 0.3 Y=1 (a) ThThe conditional distribution of Y given X is computed as shown below: P(Y =0 X =0) P(Y=0|X =0)=- P(X=0) -0.4 505 = 0.8 P(Y =1The variance of X and Y are computed as shown below: V(X)=E(X) - E(X) = 2xP(X = x)-0.52 =(0x P(X = 0))+(1+xP(X = 1)) -0.25 =0To determine the independence of X and Y verify the following expression: P(X = x, Y = y)=P(X = x)P(I = y) For different valu